OpenStudy (anonymous):

29.) Two square pyramids have the same volume. For the first pyramid, the side length of the base is 14 in, and the height is 24 in. The second pyramids has a height of 96 in. What is the side length of the base of the second pyramid? my answer: 7 in. Is that correct?

OpenStudy (anonymous):

@nincompoop @Hero @ganeshie8 @saseal

OpenStudy (anonymous):

\[\frac{ 14*w*24 }{ 3 } = \frac{ l*w*96 }{ 3 }\]

OpenStudy (anonymous):

oh wait its square...

OpenStudy (anonymous):

yea its 7

OpenStudy (anonymous):

lol thx again